Given the population: 1, 3, 4, 6, and 8; and suppose samples of size 3 are drawn from this population:






1. What is the mean and standard deviation of the population?






2. How many different samples of size n=3 can be drawn from the population? List them with their corresponding means






3. Construct the sampling distribution of the sample means.






4. What is the mean of the sampling distribution of the sample means?






5. What is the standard deviation of the sampling distribution of the sample means?

1
Expert's answer
2022-04-27T02:13:47-0400

1. We have population values 1, 3, 4, 6, 8, population size N=5 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{1+3+4+6+8}{5}=4.4"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{5}(11.56+1.96+0.16"


"+2.56+12.96)=5.84"

"\sigma=\sqrt{\sigma^2}=\sqrt{5.84}\approx2.4166"

2. The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_3=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 1,3,4 & 8/3 \\
 \hdashline
 2 & 1,3,6 & 10/3 \\
 \hdashline
 3 & 1,3,8 & 12/3 \\
 \hdashline
 4 & 1,4,6 & 11/3 \\
 \hdashline
 5 & 1,4,8 & 13/3 \\
 \hdashline
 6 & 1,6,8 & 15/3 \\
 \hdashline
 7 & 3,4,6 & 13/3 \\
 \hdashline
 8 & 3,4,8 & 15/3 \\
 \hdashline
 9 & 3,6,8 & 17/3 \\
 \hdashline
 10 & 4,6,8 & 18/3 \\
 \hdashline 
\end{array}"



3.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 8/3 & 1/10 & 8/30 & 64/90 \\
 \hdashline
 10/3 & 1/10& 10/30 & 100/90 \\
 \hdashline
 11/3 & 1/10 & 11/30 & 121/90 \\
 \hdashline
 12/3 & 1/10 & 12/30 & 144/90 \\
 \hdashline
 13/3 & 2/10 & 26/30 & 338/90 \\
 \hdashline
 15/3 & 2/10 & 30/30 & 450/90 \\
 \hdashline
 17/3 & 1/10 & 17/30 & 289/90 \\
 \hdashline
 18/3 & 1/10 & 18/30 & 324/90 \\
 \hdashline
\end{array}"



4. Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=4.4=\mu"



5. The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{1830}{90}-(4.4)^2=\dfrac{2.92}{3}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{\dfrac{2.92}{3}}\approx0.9866"